A projectile is thrown at an angle of 60 with the horizontal with an initial speed 20m/s , with H being the highest point of trajectory . Another particle is now forced to move along the same trajectory as that of the projectile such that its speed is continously increasing . When the particle P is at H ,/vp/ = 20m/s , /ap/ = 50 m/s2 the n teh acceleration vector ap at H equals.....then acceleration vector??

Grade:11

1 Answers

Sourav
6 years ago
Let the radius of curvature at pt. H be R, Now R=(v²/g)=(20cos60°)²/g=10mRadial acceleration at H= (|Vp|)²/R= 40m/s²,Tangential acceleration at H=√{ |a|² - (40)²}=30m/s²Therefore in terms of vectors net acceleration at H, (30i-40j)m/s²